# On the gcd graphs over polynomial rings and related topics

### v1

@@ 6.1. Symmetric algebras
$\psi$ is a non-degenerate $\operatorname{\mathbb{F}}_{q}$-linear functional on $\operatorname{\mathbb{F}}_{q}[x]/f.$ Consequently, $\operatorname{\mathbb{F}}_{q}[x]/f$ is a symmetric $\operatorname{\mathbb{F}}_{q}$-algebra.
@@
There exists a bijection
$\operatorname{\mathbb{F}}_{q}[x]/f\longleftrightarrow\operatorname{{\rm Hom}}((\operatorname{\mathbb{F}}_{q}[x]/f,+),\operatorname{\mathbb{C}}^{\times}),\quad a\longleftrightarrow\{\psi_{a}\},$
where
$\psi_{a}\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{C}}^{\times}$
is given by
$\psi_{a}(b)=\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ab))},\forall b\in\operatorname{\mathbb{F}}_{q}[x]/f.$
We remark that over $\operatorname{\mathbb{Z}}$, if $\zeta_{n}$ is a primitive $n$th-root of unity, then for each divisor $m\mid n$, $\zeta_{n}^{\frac{n}{m}}$ is a primitive $m$-root of unity. An analogous statement holds for $\operatorname{\mathbb{F}}_{q}[x]$ as well.
Proposition 6.10.
Let $f\in\operatorname{\mathbb{F}}_{q}[x]$ and $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}$ be a non-degenerate linear functional. Let $g$ be a divisor of $f$ and $\psi_{g}\colon\operatorname{\mathbb{F}}_{q}[x]/g\to\operatorname{\mathbb{F}}_{q}$ be the function defined by
@@
It is clear from the definition that $\psi_{g}$ is $\operatorname{\mathbb{F}}_{q}$-linear. We only need to show that it is non-degenerate. In fact, suppose to the contrary that the kernel of $\psi_{g}$ contains a non-zero ideal $I$ in $\operatorname{\mathbb{F}}_{q}[x]/g.$ Since $\operatorname{\mathbb{F}}_{q}[x]$ is a PID, $I$ must be of the form $I=\langle h\rangle$ for some $h|g.$ We then see that $\langle h\frac{f}{g}\rangle$ belongs to the kernel of $\psi.$ Because $\psi$ is non-degenerate, this implies that $h\frac{f}{g}=0$ in $\operatorname{\mathbb{F}}_{q}[x]/f.$ In other words, $g|h$ or equivalently $h=0$ in $\operatorname{\mathbb{F}}_{q}[x]/g.$ This shows that $I=0$, which is a contradiction. ∎
@@ 6.2. Ramanujan sums over $\operatorname{\mathbb{F}}_{q}[x]$
(Ramanujan sums over $\operatorname{\mathbb{F}}_{q}[x]$) Let $f,g\in\operatorname{\mathbb{F}}_{q}[x]$ be monic polynomials. Let $\psi:\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}$ be a non-degenerate linear functional on $\operatorname{\mathbb{F}}_{q}[x]/f$. The Ramanujan sum $c(g,f)$ is defined as
$c(g,f)=c_{\psi}(g,f)=\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ga))}.$
@@
We remark also that at first glance, $c(g,f)$ depends on the choice of $\psi.$ However, as we explain in what follows, it does not as long as we make sure that $\psi$ is non-degenerate. This is similar to the case over $\operatorname{\mathbb{Z}}$: Ramanujan sums do not depend on the choice of a primitive $n$th-root of unity. In fact, we will show that there is an explicit formula for $c(g,f)$ similar to the case of Ramanujan sum over $\operatorname{\mathbb{Z}}.$
@@
Ramanujan sums are a special case of Gauss sums as defined and studied in [21, Definition 1]. In fact, they are Gauss sums for the principal Dirichlet characters on $\operatorname{\mathbb{F}}_{q}[x]/f.$ It would be rather interesting if we could define Fekete polynomials for these principal Dirichlet characters (the case over $\operatorname{\mathbb{Z}}$ was studied in [7]).
@@
Let $\psi$ be a non-degenerate linear functional on $\operatorname{\mathbb{F}}_{q}[x]/f$. For each $f\in\operatorname{\mathbb{F}}_{q}[x]$
@@
If $f=\prod_{i=1}^{d}f_{i}^{\alpha_{i}}$ where $\alpha_{i}\in\operatorname{\mathbb{N}}$ and $f_{i}$ are irreducible, then by the Chinese remainder theorem and [21, Satz 1] we have
@@
where $\psi_{i}:\operatorname{\mathbb{F}}_{q}[x]/f_{i}^{\alpha_{i}}\to\operatorname{\mathbb{F}}_{q}$ is a non-degenerate linear functional.
@@
$\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/f_{1}^{a_{1}-1}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(f_{1}a))}=0.$
On the other hand if $a_{1}=1$ then $\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/f_{1}^{a_{1}-1}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(f_{1}a))}=-1$. We conclude that
@@
Let $f,g\in\operatorname{\mathbb{F}}_{q}[x]$ be monic polynomials. Then (compare with [18, Equation 9])
@@
Let $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/f\to\operatorname{\mathbb{F}}_{q}$ be a non-degenerate linear functional and
@@
Let $h=\gcd(g,f)$ and $t=\dfrac{f}{h}$. The canonical projection map $(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}\to(\operatorname{\mathbb{F}}_{q}[x]/t)^{\times}$ has kernel of size $\frac{\varphi(f)}{\varphi(t)}.$ Furthermore, for each $a\in(\operatorname{\mathbb{F}}_{q}[x]/(t))^{\times}$, if $a_{1},a_{2}$ are two preimages of $a$ in $(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}$ then $a_{2}=a_{1}+tu$, for some $u\in\operatorname{\mathbb{F}}_{q}[x]/f$. In this case, one has $\psi(ga_{2})=\psi(ag_{1}+gtu)=\psi(ga_{1})+\psi(\frac{gu}{h}f)=\psi(ga_{1})$. Consequently, we have
$c_{\psi}(g,f)$ $=\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/f)^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ga))}=\frac{\varphi(f)}{\varphi(t)}\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/t)^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ga))}$
@@
Here $\psi_{t}:\operatorname{\mathbb{F}}_{q}[x]/t\to\operatorname{\mathbb{F}}_{q}$ is the linear functional defined by $\psi_{t}(b)=\psi(\frac{f}{t}b)=\psi(hb)$. Note that $\psi_{t}$ is non-degenerate.
@@ 6.3. Spectrum of $G_{f}(D)$
Let $G$ be an abelian group and $S$ a symmetric subset of $G$ such that $0\not\in S$. The adjacency matrix of the Cayley graph $\Gamma(G,S)$ is a $G$-circulant matrix. By the $G$-circulant theorem, the spectrum of $\Gamma(G,S)$ is the collection of the sums $\sum_{s\in S}\chi(s)$, where $\chi$ runs over the set of all characters of $G$ (see [14, Section 1.2]). When $G=\operatorname{\mathbb{F}}_{q}[x]/f$, we know by 6.9 that the character of $G$ is parameterized by $G$ itself. Consequently, we have the following proposition.
@@
Let $S$ be a subset of $\operatorname{\mathbb{F}}_{q}[x].$ Then the spectrum of the Cayley graph $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/f,S)$ is given by the set
$\left\{\sum_{s\in S}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))}\right\}_{g\in\operatorname{\mathbb{F}}_{q}[x]/f}.$
@@
$\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/f,\gcd(a,f)=h}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ag))}=c\left(g,\frac{f}{h}\right).$
@@
$\sum_{b\in\operatorname{\mathbb{F}}_{q}[x]/(f/h),\gcd(a,f/h)=1}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(bgh))}=\sum_{b\in\operatorname{\mathbb{F}}_{q}[x]/(f/h),\gcd(b,f/h)=1}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi_{h}(bg))}.$
Here $\psi_{h}:\operatorname{\mathbb{F}}_{q}[x]/(f/h)\to\operatorname{\mathbb{F}}_{q}$ is the functional given by $\psi_{h}(x)=\psi(hx).$ By 6.10, $\psi_{h}$ is a non-generate linear functional on $\operatorname{\mathbb{F}}_{q}[x]/(f/g).$ Therefore, we conclude that
$\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/f,\gcd(a,f)=h}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ag))}=\sum_{b\in\operatorname{\mathbb{F}}_{q}[x]/(f/h),\gcd(b,f/h)=1}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi_{h}(bg))}=c\left(g,\frac{f}{h}\right).$
@@
Let $f$ be a monic polynomial and $D=\{f_{1},f_{2},\ldots,f_{k}\}$ where $f_{i}\mid f.$ Then, the spectrum of $S_{f}(D)$ is given by the set
$\left\{\sum_{i=1}^{k}c\left(g,\frac{f}{f_{i}}\right)\right\}_{g\in\operatorname{\mathbb{F}}_{q}[x]/f}.$
Corollary 6.20.
All eigenvalues of the gcd-graph $G_{f}(D)$ are integers.
As we discussed in the introduction, it is known that a $\operatorname{\mathbb{Z}}/n$-circulant graph has an integral spectrum if and only if it is a gcd-graph (see [29, Theorem 7.1] ). One may ask whether the same statement is true for graphs associated with $\operatorname{\mathbb{F}}_{q}[x]/f.$ The answer is no in general. In fact, we have the following general observation.
@@
Let $S$ be a symmetric subset of $\operatorname{\mathbb{F}}_{q}[x]/f$ such that $0\not\in S.$ Suppose further that $\operatorname{\mathbb{F}}_{p}^{\times}S=S.$ Then, the Cayley graph $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/f,S)$ has an integral spectrum.
@@
We define the following equivalence relation on $\operatorname{\mathbb{F}}_{q}[x]/f$. For $u,v\in\operatorname{\mathbb{F}}_{q}[x]/f$, we say that $u\sim v$ if $u=kv$ where $k\in\operatorname{\mathbb{F}}_{p}^{\times}.$ By 6.17, it is enough to show that for each $g\in\operatorname{\mathbb{F}}_{q}[x]/f$, $\sum_{s\in S}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))}\in\operatorname{\mathbb{Z}}$. Because $\operatorname{\mathbb{F}}_{p}^{\times}S=S$, this sum can be written as
@@
This shows that $\sum_{s\in S}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))}\in\operatorname{\mathbb{Z}}$ for each $g\in\operatorname{\mathbb{F}}_{q}[x]/f.$
@@
Let $\operatorname{\mathbb{F}}_{q}$ be a finite field such that $\operatorname{\mathbb{F}}_{q}\neq\operatorname{\mathbb{F}}_{p}.$ Let $f=x$. In this case $\operatorname{\mathbb{F}}_{q}[x]/f\cong\operatorname{\mathbb{F}}_{q}.$ Let $V$ be proper $\operatorname{\mathbb{F}}_{p}$-subspace of $\operatorname{\mathbb{F}}_{q}$ and $S=V\setminus\{0\}.$ Then by 6.21, $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/f,S)$ has an integral spectrum even though it is not a gcd-graph.
We wonder whether the converse of 6.21 holds (perhaps, under some mild assumptions).

### v2

@@ 6.1. Symmetric algebras
$\psi$ is a non-degenerate $\operatorname{\mathbb{F}}_{q}$-linear functional on $\operatorname{\mathbb{F}}_{q}[x]/(f).$ Consequently, $\operatorname{\mathbb{F}}_{q}[x]/(f)$ is a symmetric $\operatorname{\mathbb{F}}_{q}$-algebra.
By 6.4, under the composition $\operatorname{\mathbb{F}}_{q}[x]/(f)\xrightarrow{\psi}\operatorname{\mathbb{F}}_{q}\xrightarrow{\operatorname{{\rm Tr}}}\operatorname{\mathbb{F}}_{p}$ where Tr is the trace map, $\operatorname{\mathbb{F}}_{q}[x]/(f)$ becomes a symmetric $\operatorname{\mathbb{F}}_{p}$-algebra. By 6.6, we have the following proposition.
Proposition 6.9.
There exists a bijection
$\operatorname{\mathbb{F}}_{q}[x]/(f)\longleftrightarrow\operatorname{{\rm Hom}}((\operatorname{\mathbb{F}}_{q}[x]/(f),+),\operatorname{\mathbb{C}}^{\times}),\quad a\longleftrightarrow\{\psi_{a}\},$
where
$\psi_{a}\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{C}}^{\times}$
is given by
$\psi_{a}(b)=\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ab))},\forall b\in\operatorname{\mathbb{F}}_{q}[x]/(f).$
@@
Let $f\in\operatorname{\mathbb{F}}_{q}[x]$ and $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}$ be a non-degenerate linear functional. Let $g$ be a divisor of $f$ and $\psi_{g}\colon\operatorname{\mathbb{F}}_{q}[x]/g\to\operatorname{\mathbb{F}}_{q}$ be the function defined by
@@
It is clear from the definition that $\psi_{g}$ is $\operatorname{\mathbb{F}}_{q}$-linear. We only need to show that it is non-degenerate. In fact, suppose to the contrary that the kernel of $\psi_{g}$ contains a non-zero ideal $I$ in $\operatorname{\mathbb{F}}_{q}[x]/g.$ Since $\operatorname{\mathbb{F}}_{q}[x]$ is a PID, $I$ must be of the form $I=\langle h\rangle$ for some $h\mid g.$ We then see that $\langle h\frac{f}{g}\rangle$ belongs to the kernel of $\psi.$ Because $\psi$ is non-degenerate, this implies that $h\frac{f}{g}=0$ in $\operatorname{\mathbb{F}}_{q}[x]/(f).$ In other words, $g\mid h$ or equivalently $h=0$ in $\operatorname{\mathbb{F}}_{q}[x]/g.$ This shows that $I=0$, which is a contradiction. ∎
@@ 6.2. Ramanujan sums over $\operatorname{\mathbb{F}}_{q}[x]$
(Ramanujan sums over $\operatorname{\mathbb{F}}_{q}[x]$) Let $f,g\in\operatorname{\mathbb{F}}_{q}[x]$ be polynomials such that $f$ is monic. Let $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}$ be a non-degenerate linear functional on $\operatorname{\mathbb{F}}_{q}[x]/(f)$. The Ramanujan sum $c(g,f)$ is defined as
$c(g,f)=c_{\psi}(g,f)=\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/(f))^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ga))}.$
@@
[…]
@@
[…]
@@
The Möbius function $\mu$ on $\operatorname{\mathbb{F}}_{q}[x]$ is defined in the same way as in the integer case, which we recall as follows: for a non-zero polynomial $f\in\operatorname{\mathbb{F}}_{q}[x]$,
$\mu(f)=\begin{cases}(-1)^{r}&\text{if $f=aP_{1}\cdots P_{r}$, }\\ 0&\text{ if $f$ is not square-free.}\end{cases}$
Here $a$ is a non-zero constant and $P_{1},\ldots,P_{r}$ are distinct monic irreducible polynomials dividing $f$ ($r$ could be 0).
@@
Let $\psi$ be a non-degenerate linear functional on $\operatorname{\mathbb{F}}_{q}[x]/(f)$. For each $f\in\operatorname{\mathbb{F}}_{q}[x]$
@@
If $f=\prod_{i=1}^{d}f_{i}^{\alpha_{i}}$ where $\alpha_{i}\in\operatorname{\mathbb{N}}$ and $f_{i}$ are irreducible, then by the Chinese remainder theorem and [22, Satz 1] we have
@@
where $\psi_{i}\colon\operatorname{\mathbb{F}}_{q}[x]/(f)_{i}^{\alpha_{i}}\to\operatorname{\mathbb{F}}_{q}$ is a non-degenerate linear functional induced by $\psi.$
@@
$\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/(f_{1}^{a_{1}-1})}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(f_{1}a))}=0.$
On the other hand if $a_{1}=1$ then $\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/(f_{1}^{a_{1}-1})}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(f_{1}a))}=1$. We conclude that
@@
Let $f,g\in\operatorname{\mathbb{F}}_{q}[x]$ be polynomials such that $f$ is monic. Then (compare with [19, Equation 9])
@@
Here $\varphi$ is the Euler totient function on $\operatorname{\mathbb{F}}_{q}[x]$; i.e, $\varphi(h)=|(\operatorname{\mathbb{F}}_{q}[x]/(h))^{\times}|.$ In particular, $c(g,f)$ is an integer.
@@
Let $\psi\colon\operatorname{\mathbb{F}}_{q}[x]/(f)\to\operatorname{\mathbb{F}}_{q}$ be a non-degenerate linear functional and
@@
Let $h=\gcd(g,f)$ and $t=\dfrac{f}{h}$. The canonical projection map
$(\operatorname{\mathbb{F}}_{q}[x]/(f))^{\times}\to(\operatorname{\mathbb{F}}_{q}[x]/(t))^{\times}$},
has kernel of size $\frac{\varphi(f)}{\varphi(t)}.$ Furthermore, for each $a\in(\operatorname{\mathbb{F}}_{q}[x]/(t))^{\times}$, if $a_{1},a_{2}$ are two preimages of $a$ in $(\operatorname{\mathbb{F}}_{q}[x]/(f))^{\times}$ then $a_{2}=a_{1}+tu$, for some $u\in\operatorname{\mathbb{F}}_{q}[x]/(f)$. In this case, one has
$\psi(ga_{2})=\psi(ag_{1}+gtu)=\psi(ga_{1})+\psi(\frac{gu}{h}f)=\psi(ga_{1}).$
Consequently, we have
$c_{\psi}(g,f)$ $=\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/(f))^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ga))}=\frac{\varphi(f)}{\varphi(t)}\sum_{a\in(\operatorname{\mathbb{F}}_{q}[x]/(t))^{\times}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ga))}$
@@
Here $\psi_{t}\colon\operatorname{\mathbb{F}}_{q}[x]/(t)\to\operatorname{\mathbb{F}}_{q}$ is the linear functional defined by $\psi_{t}(b)=\psi(\frac{f}{t}b)=\psi(hb)$. Note that $\psi_{t}$ is non-degenerate.
@@ 6.3. Spectrum of $G_{f}(D)$
Let $G$ be an abelian group and $S$ a symmetric subset of $G$ such that $0\not\in S$. The adjacency matrix of the Cayley graph $\Gamma(G,S)$ is a $G$-circulant matrix. By the $G$-circulant theorem, the spectrum of $\Gamma(G,S)$ is the collection of the sums $\sum_{s\in S}\chi(s)$, where $\chi$ runs over the set of all characters of $G$ (see [15, Section 1.2, Equation (1.1’)]). When $G=\operatorname{\mathbb{F}}_{q}[x]/(f)$, we know by 6.9 that the character of $G$ is parameterized by $G$ itself. Consequently, we have the following proposition.
@@
Let $S$ be a subset of $\operatorname{\mathbb{F}}_{q}[x].$ Then the eigenvalues of the Cayley graph $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/(f),S)$ are given by the set
$\left\{\sum_{s\in S}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))}\right\}_{g\in\operatorname{\mathbb{F}}_{q}[x]/(f)}.$
@@
$\sum_{a\in\operatorname{\mathbb{F}}_{q}[x]/(f),\gcd(a,f)=h}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ag))}=c\left(g,\frac{f}{h}\right).$
@@
$\sum_{\begin{subarray}{c}b\in\operatorname{\mathbb{F}}_{q}[x]/(f/h)\\ \gcd(a,f/h)=1\end{subarray}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(bgh))}=\sum_{\begin{subarray}{c}b\in\operatorname{\mathbb{F}}_{q}[x]/(f/h)\\ \gcd(b,f/h)=1\end{subarray}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi_{h}(bg))}.$
Here $\psi_{h}\colon\operatorname{\mathbb{F}}_{q}[x]/(f/h)\to\operatorname{\mathbb{F}}_{q}$ is the functional given by $\psi_{h}(x)=\psi(hx).$ By 6.10, $\psi_{h}$ is a non-generate linear functional on $\operatorname{\mathbb{F}}_{q}[x]/(f/g).$ Therefore, we conclude that
$\sum_{\begin{subarray}{c}a\in\operatorname{\mathbb{F}}_{q}[x]/(f)\\ \gcd(a,f)=h\end{subarray}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(ag))}=\sum_{\begin{subarray}{c}b\in\operatorname{\mathbb{F}}_{q}[x]/(f/h)\\ \gcd(b,f/h)=1\end{subarray}}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi_{h}(bg))}=c\left(g,\frac{f}{h}\right).$
@@
Let $f$ be a monic polynomial and $D=\{f_{1},f_{2},\ldots,f_{k}\}$ where $f_{i}\mid f.$ Then, the eigenvalues of $G_{f}(D)$ are given by the set
$\left\{\sum_{i=1}^{k}c\left(g,\frac{f}{f_{i}}\right)\right\}_{g\in\operatorname{\mathbb{F}}_{q}[x]/(f)}.$
In particular, all eigenvalues of the gcd-graph $G_{f}(D)$ are integers.
Example 6.20.
The characteristic polynomial of the graph described in Fig. 1 has the following factorization
$(x-4)(x-1)^{4}(x+2)^{4}.$
Similarly, the characteristic polynomial of the graph described in Fig. 2 has the following factorization
$(x-3)(x+3)(x-1)^{3}(x+1)^{3}.$
In particular, the eigenvalues of these graphs are integers as shown in 6.19. We remark that these spectral calculations have also been verified using Ramanujan sums, as explained in 6.19. These verifications are described in [24], which contains the code to calculate various arithmetic functions over $\operatorname{\mathbb{F}}_{q}[x]$. In particular, it has a function to calculate the general Ramanujan sum $c(g,f)$ for all $g,f\in\operatorname{\mathbb{F}}_{q}[x].$
For example, let us consider the graph $G_{f}(D)$ described in Fig. 1, where $f=x(x+1)\in\operatorname{\mathbb{F}}_{3}[x]$} and $D=\{x,x+1\}$. We shall identify $\operatorname{\mathbb{F}}_{3}[x]/(f)$ with
$\{0,1,2,x,x+1,x+2,2x,2x+1,2x+2\}.$
For $g=0$, we have $c(0,x)=\mu(1)\dfrac{\varphi(x)}{\varphi(1)}=2$, and similarly $c(0,x+1)=2$. Hence the eigenvalue of $G_{f}(D)$ associated with $g=0$ is $c(0,x)+c(0,x+1)=4$. (The eigenvalue $4$ is the degree of $G_{f}(D)$, and it appears with multiplicity $1$.)
For $g$ coprime to $x$, we have $c(g,x)=\mu(x)\dfrac{\varphi(x)}{\varphi(x)}=-1$. Similarly, for $g$ coprime to $x+1$, we have c(g,x+1)=1c(g,x+1)=-1. We also have
$c(x,x)=c(2x,x)=\mu(1)\dfrac{\varphi(x)}{\varphi(1)}=2,$
and
$c(x+1,x+1)=c(2x+2,x+1)=2.$
From the above discussion, we see that if $g\in\{1,2,x+2,2x+1\}$, then the eigenvalue of $G_{f}(D)$ associated with $g$ is
c(g,x)+c(g,x+1)=(1)+(1)=2.c(g,x)+c(g,x+1)=(-1)+(-1)=-2.
(The eigenvalue 2-2 appears with multiplicity 4.) Finally, if $g\in\{x,x+1,2x,2x+2\}$, then the eigenvalue of $G_{f}(D)$ associated with $g$ is
c(g,x)+c(g,x+1)=2+(1)=1.c(g,x)+c(g,x+1)=2+(-1)=1.
(The eigenvalue $1$ appears with multiplicity 4.) Our calculations show that the spectrum of $G_{f}(D)$, as described in 6.19, coincides with the spectrum obtained from the adjacency matrix.
As we discussed in the introduction, it is known that a $\operatorname{\mathbb{Z}}/n\operatorname{\mathbb{Z}}$-circulant graph has an integral spectrum if and only if it is a gcd-graph (see [32, Theorem 7.1] ). One may ask whether the same statement is true for graphs associated with $\operatorname{\mathbb{F}}_{q}[x]/(f).$ The answer is no in general. In fact, we have the following general observation.
@@
Let $S$ be a symmetric subset of $\operatorname{\mathbb{F}}_{q}[x]/(f)$ such that $0\not\in S.$ Suppose further that $\operatorname{\mathbb{F}}_{p}^{\times}S=S.$ Then, the Cayley graph $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/(f),S)$ has an integral spectrum.
@@
We define the following equivalence relation on $\operatorname{\mathbb{F}}_{q}[x]/(f)$. For $u,v\in\operatorname{\mathbb{F}}_{q}[x]/(f)$, we say that $u\sim v$ if $u=kv$ where $k\in\operatorname{\mathbb{F}}_{p}^{\times}.$ By 6.17, it is enough to show that for each $g\in\operatorname{\mathbb{F}}_{q}[x]/(f)$, $\sum_{s\in S}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))}\in\operatorname{\mathbb{Z}}$. Because $\operatorname{\mathbb{F}}_{p}^{\times}S=S$, this sum can be written as
@@
This shows that $\sum_{s\in S}\zeta_{p}^{\operatorname{{\rm Tr}}(\psi(gs))}\in\operatorname{\mathbb{Z}}$ for each $g\in\operatorname{\mathbb{F}}_{q}[x]/(f).$
@@
Let $\operatorname{\mathbb{F}}_{q}$ be a finite field such that $\operatorname{\mathbb{F}}_{q}\neq\operatorname{\mathbb{F}}_{p}.$ Let $f=x$. In this case $\operatorname{\mathbb{F}}_{q}[x]/(f)\cong\operatorname{\mathbb{F}}_{q}.$ Let $V$ be proper $\operatorname{\mathbb{F}}_{p}$-subspace of $\operatorname{\mathbb{F}}_{q}$ and $S=V\setminus\{0\}.$ Then by 6.21, $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/(f),S)$ has an integral spectrum even though it is not a gcd-graph. In a recent work, we show that the converse of this statement is also true; namely, $\Gamma(\operatorname{\mathbb{F}}_{q}[x]/(f),S)$ is integral if and only if $S\cup\{0\}$ is a vector space over $\operatorname{\mathbb{F}}_{p}$ (see [28, Theorem 1.1].)